# Toward Unified Practices in Trajectory Prediction Research on Drone Datasets

### v3

@@ Abstract
The availability of high-quality datasets is crucial for the development of behavior prediction algorithms in autonomous vehicles. This paper highlights the need to standardize the use of certain datasets for motion forecasting research to simplify comparative analysis and proposes a set of tools and practices to achieve this. Drawing on extensive experience and a comprehensive review of current literature, we summarize our proposals for preprocessing, visualization, and evaluation in the form of an open-sourced toolbox designed for researchers working on trajectory prediction problems. The clear specification of necessary preprocessing steps and evaluation metrics is intended to alleviate development efforts and facilitate the comparison of results across different studies. The toolbox is available at: https://github.com/westny/dronalize.
@@ I Introduction
For a long time, researchers have relied on traffic datasets such as the I-80, US-101, Lankershim, and Peachtree datasets from the Next Generation Simulation (NGSIM) program [3] that have been widely used for studies on behavior prediction problems [4, 5, 6, 7, 8, 9, 10]. The NGSIM datasets include data from the perspective of a stationary camera overlooking a region of interest, providing a birds-eye view (BEV) of the traffic scene. However, some of these datasets suffer from high noise levels and tracking errors, which makes them less reliable for traffic studies [11, 12]. In response, recent years have seen the release of new datasets that aim to address some of these issues, such as SDD [13], INTERACTION [14], SIND [15] and those from leveLXData [16, 17, 18, 19, 20]. These datasets are primarily collected using camera-equipped aerial drones and provide data of particularly high spatial and temporal resolution. However, a challenge in using some of these datasets is the lack of a widely recognized standard for preprocessing steps—making it difficult to reproduce results and compare performances across various trajectory prediction studies.
In contrast to the aforementioned datasets, there are also those collected by intelligent vehicles, e.g., HDD [21], NuScenes [22], Argoverse 2 [23], and Waymo Open Motion Dataset (WOMD) [24], which consist of data from onboard sensors, and thus from the point of view (POV) of the vehicle. These datasets have seen widespread adoption in the research community, in particular Argoverse 2 and WOMD, as they provide benchmarking suites for trajectory prediction problems and appear in workshop challenges [25]. Because the data are collected using the actual sensor suite of the vehicle, their use facilitates model development adopted for real-world conditions. Despite their realism, however, these datasets may not always capture all surrounding objects that influence the behavior of other agents because of potential sensor occlusions. This aspect introduces uncertainty in evaluations, making it challenging to discern the specific contributions of the model and the data in highly interactive scenarios. Importantly, we argue that neither BEV or POV-style datasets are preferable over the other; rather, they should be complementary in trajectory prediction studies. Inspired by the most recent POV-style datasets, we introduce a standardized preprocessing pipeline and evaluation protocol for trajectory prediction research on several drone datasets; in particular, highD [16], rounD [17], inD [18], exiD [19], uniD [20], SIND [15], A43 [26], and INTERACTION [14]. By reconciling various approaches from the existing literature, we aim to establish a standard that reflects the most common practices in the field.
@@ I-A Contributions
tion of this paper is the proposal of standardized methodologies for trajectory prediction research on drone datasets. These practices are realized in the form of a toolbox that serves two primary purposes:
@@
The focus on the aforementioned datasets is driven by several factors. First, their high tracking accuracy allows for interpretable and reliable evaluations—making them particularly suitable for early-stage research and development. Secondly, these datasets feature a diverse range of interacting road users, including cars, trucks, and vulnerable road users (VRUs) such as pedestrians and cyclists, providing rich scenarios for studying safety-critical interactions in autonomous driving applications. The data were collected at several fixed locations (highways, intersections, roundabouts, etc.) across Germany, China, and the United States, providing a broad range of traffic scenarios. The datasets are made up of multiple recordings that contain both static and dynamic information, provided in .csv files. All but highD are also accompanied by a Lanelet 2 [28] file in .osm format, which can be used to extract semantic information about the scene.
II-A Related Work
Although there are numerous examples of studies that utilize the aforementioned datasets for behavior prediction research [29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43], they are used in different ways, some of which is illustrated in Table I. These inconsistencies complicate the comparison of results across studies since making fair comparisons necessitates uniform conditions for all methods. Rather than advocating for one approach over another, we aim to propose a standard that reflects the most common practices in the field. As such, there is a connection between this work and the ones that organize challenges or workshops, e.g., [22], [23], or [24], as they provide a clear specification of dataset partitioning and evaluation metrics.
TABLE I: Time specifications of studies using drone datasets.
@@ III-B Dataset Splits
It is important to recognize that there is a notable imbalance across the maneuver classes within the highway datasets (the lane-change percentage is around $13\%$). To ensure that the scenarios are both dynamic and challenging, we adopt a stratified sampling approach, aiming to achieve a more balanced distribution of the TA maneuvers—the results of which are shown in Table III. We use this information to label the trajectories with the corresponding maneuver class, which can be used for assessing the model’s ability to predict specific maneuvers as well as intention prediction applications. Seven maneuver classes are employed based on the time-to-lane-change (TTLC), as described in Fig. 3.
@@ III-C Coordinate System
Although the datasets are consistent in terms of the data they provide, some differences need to be addressed, such as the coordinate system used. This is mainly an issue when working with the highD dataset. To ensure that the samples are consistent across the datasets, the data is first divided into two separate groups based on the direction of travel. Next, the data are transformed to a common coordinate system where the $x$-axis points in the direction of the road, the $y$-axis points to the left, and the origin is moved to the lower left corner of the road. For the other datasets, there are no such obvious issues, as all agents move in the same area. Regardless, we still place the origin in the approximate geographical center of the scene to ensure some consistency across recordings.
@@
A potential issue with using a global coordinate system is that it may introduce prediction bias, where the model learns scene-specific patterns rather than generalizable behavior—an essential property for zero-shot learning [44, 45]. While we leave this design choice to the user, one possible approach is to transform the scenario into a relative coordinate system, such as the local frame of the TA.
III-D Downsampling
@@
Most related works have used downsampling as a preprocessing step [29, 37, 34, 33, 35, 46, 38, 36, 39, 40, 41, 43], albeit with different integer factors. Although there are several potential computational benefits—allowing for faster prototyping and experimentation—it is important to consider the qualitative effect downsampling has on both the data and the model. Figure 4 is used to illustrate how different sampling rates can affect the quality of predicted trajectories using nominal driving conditions for the rounD dataset using a 1st-order method. In the figure, we see that for lower sampling rates, the simulated trajectories become overly jagged and unrealistic, which makes it difficult to assess the physical feasibility of the predicted trajectories. To ensure that the data are still representative of the original tracking data, we propose to downsample it to a rate of $5$ Hz, corresponding to a sampling time of $0.2$ s. When studying the frequency spectrum of the tracking data, it was found that several signals have components above the new Nyquist frequency ($2.5$ Hz). Before downsampling, an anti-aliasing filter is applied to remove these high-frequency components using a 7th order Chebyshev type I filter (based on qualitative evaluations).
@@
Figure 4: Example simulation of a vehicle moving along a circular path with a radius of $16$ m, a speed of $50$ km/h (speed limit) for a duration of $5$ s using different sampling rates in a forward Euler integration scheme. The values represent nominal conditions for the rounD dataset. The figure is used to illustrate how different sampling rates can affect motion prediction quality. Something that could be of interest when assessing the physical feasibility of the predicted trajectories.
III-E Agent Features
III-E1 Tracking Features
@@
The longitudinal and lateral accelerations are included as supplementary features, i.e., for development purposes. We argue against their inclusion in the final model since acceleration measurements are rarely available in POV data—they could of course be estimated online, e.g., using finite differences or serve a purpose in an extended ablation study.
@@
III-E2 Agent Classes
The different datasets contain agents of different types, e.g., cars, bicycles, pedestrians, etc. We hypothesize that the model could benefit from the inclusion of this information, as different classes exhibit different behaviors. To allow for a more generalizable model, we propose a common set of agent classes across the datasets. The categories could be included as one-hot encoded features, or directly used as tokens in an embedding layer. The distribution of unique agents and their respective classes in the datasets is shown in Table V. We note that the number of unique agents in the INTERACTION dataset is not explicitly provided. However, the dataset paper reports a total of 40,054 recorded vehicles. For reference, our preprocessed data contains 676,931 vehicle trajectories, along with 32,722 trajectories corresponding to a single VRU class.
@@ III-E2 Agent Classes
Figure 5: Example lane graphs for the rounD, inD, SIND, and INTERACTION datasets. The lane graphs are constructed using the available Lanelet files, which contain semantic information about the scene. The lane graphs are constructed by extracting lane boundaries and markings, coloring them according to their type, and connecting them with edges. Although these examples are simple, combined with the available traffic data, they could be sufficient for many trajectory prediction tasks. Using the provided functionality, users can develop more complex lane graph constructions based on the desired detail level.
III-F High-Definition Maps
@@
All but the highD dataset are accompanied by Lanelet 2 [28] files that can be used to extract semantic information about the scenes. Our toolbox includes functionality for basic lane-graph construction, designed to serve as a foundation that users can modify according to their specific needs and preferences. Although the highD dataset does not include Lanelets, available geographical information can be used to construct a simple lane graph. Examples are shown in Fig. 5.
III-G Data Structure
@@
IV Evaluation Metrics
This section outlines the most commonly used metrics from the literature to evaluate motion prediction models, which are incorporated in the toolbox. The metrics are here presented for a single agent but can easily be extended to multi-agent problems by averaging over all scored agents in the scenario. Here, $\bm{x}_{k}\in\mathbb{R}^{2}$ refers to the ground truth position at time step $k$, $\hat{\bm{x}}_{k}$ refers to the respective prediction, and $N$ is the prediction horizon steps.
@@ IV Evaluation Metrics
V Future Directions
The introduction of the toolbox for the development of trajectory-prediction models for drone datasets opens up several directions for future research. Here we outline some potential areas for further exploration.
@@ V Future Directions
  • Generalization & Zero-Shot Learning: Explore model generalization and zero-shot learning across datasets to investigate how models can be applied or adapted across different scenarios.
  • Transfer Learning: Explore the potential of applying models, initially trained on one or more drone datasets with subsequent fine-tuning, to other datasets within the suite or even to external datasets such as NuScenes, Argoverse 2, or WOMD. An example of how transfer learning can be applied to the INTERACTION and inD datasets using a simple encoder-decoder model is shown in Fig. 6.
Figure 6: Illustration of the impact of transfer learning on learning performance for the inD and INTERACTION datasets. The figure shows how validation error evolves over training epochs, where fine-tuning a model on one dataset after pretraining on the other leads to a lower initial error that remains consistently lower throughout training. These results highlight the potential of transfer learning for improved generalization and motivate further research in this direction.
VI Conclusions
This paper has presented a set of methodologies for the development of trajectory prediction models using drone datasets. The proposals have been motivated by an analysis of the most common practices in field. The methods have been implemented in an open-source toolbox designed to standardize preprocessing and performance evaluation procedures—thereby simplifying comparative studies. By alleviating the technical burden of developing research pipelines, the toolbox augments the potential for exploration and discoveries in the field of trajectory prediction.

### v4

# Toward Unified Practices in Trajectory Prediction Research on Bird’s-Eye-View Datasets

@@ Abstract
The availability of high-quality datasets is crucial for developing behavior prediction algorithms in autonomous vehicles. This paper highlights the need to standardize the use of certain datasets for motion forecasting research to simplify comparative analysis and proposes a set of tools and practices to achieve this. Drawing on extensive experience and a comprehensive review of current literature, we summarize our proposals for preprocessing, visualization, and evaluation in the form of an open-sourced toolbox designed for researchers working on trajectory prediction problems. The clear specification of necessary preprocessing steps and evaluation metrics is intended to alleviate development efforts and facilitate the comparison of results across different studies. The toolbox is available at: https://github.com/westny/dronalize.
@@ I Introduction
For a long time, researchers have relied on traffic datasets such as the I-80, US-101, Lankershim, and Peachtree datasets from the Next Generation Simulation (NGSIM) program [3] that have been widely used for studies on behavior prediction problems [4, 5, 6, 7, 8, 9, 10]. The NGSIM datasets include data from the perspective of a stationary camera overlooking a region of interest, providing a bird’s-eye view (BEV) of the traffic scene. However, some of these datasets suffer from high noise levels and tracking errors, which makes them less reliable for traffic studies [11, 12]. In response, recent years have seen the release of new datasets that aim to address some of these issues, such as SDD [13], INTERACTION [14], SIND [15] and those from leveLXData [16, 17, 18, 19, 20]. These datasets are primarily collected using camera-equipped aerial drones and provide data of particularly high spatial and temporal resolution. However, a challenge in using some of these datasets is the lack of a widely recognized standard for preprocessing steps—making it difficult to reproduce results and compare performances across various trajectory prediction studies.
In contrast to the aforementioned datasets, there are also those collected by intelligent vehicles, e.g., HDD [21], NuScenes [22], Argoverse 2 [23], and Waymo Open Motion Dataset (WOMD) [24], which consist of data from onboard sensors, and thus from the point of view (POV) of the vehicle. These datasets have seen widespread adoption in the research community, in particular Argoverse 2 and WOMD, as they provide benchmarking suites for trajectory prediction problems and appear in workshop challenges [25]. Because the data are collected using the actual sensor suite of the vehicle, their use facilitates model development adopted for real-world conditions. Despite their realism, however, these datasets may not always capture all surrounding objects that influence the behavior of other agents because of potential sensor occlusions. This aspect introduces uncertainty in evaluations, making it challenging to discern the specific contributions of the model and the data in highly interactive scenarios. Importantly, we argue that neither BEV nor POV type of datasets are preferable over the other; rather, they should be complementary in trajectory prediction studies. Inspired by the most recent POV-style datasets, we introduce a standardized preprocessing pipeline and evaluation protocol for trajectory prediction research on several BEV datasets; in particular, highD [16], rounD [17], inD [18], exiD [19], uniD [20], SIND [15], A43 [26], and INTERACTION [14]. By reconciling various approaches from the existing literature, we aim to establish a standard that reflects the most common practices in the field.
@@ I-A Contributions
The contribution of this paper is the proposal of standardized practices for trajectory prediction research on BEV datasets. These practices are realized in the form of a toolbox that serves two primary purposes:
@@ II Background
The focus on the aforementioned datasets is driven by several factors. First, their high tracking accuracy allows for interpretable and reliable evaluations—making them particularly suitable for early-stage research and development. Second, these datasets feature a diverse range of interacting road users, including cars, trucks, and vulnerable road users (VRUs) such as pedestrians and cyclists, providing rich scenarios for studying safety-critical interactions in autonomous driving applications. The data were collected at several fixed locations (highways, intersections, roundabouts, etc.) across Germany, China, and the United States, providing a broad range of traffic scenarios. The datasets are made up of multiple recordings that contain both static and dynamic information, provided in .csv files. All but highD are also accompanied by a Lanelet 2 [28] file in .osm format, which can be used to extract semantic information about the scene.
III Related Work
Although there are numerous examples of studies that utilize the aforementioned datasets for behavior prediction research [29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43], they are used in different ways, some of which are illustrated in Table I. These inconsistencies complicate the comparison of results across studies, as fair evaluation requires uniform conditions for all methods. Rather than advocating for one approach over another, we aim to propose a standard that reflects the most common practices in the field. Related efforts include tools that provide unified interfaces to facilitate data access and evaluation across well-curated trajectory prediction datasets, such as [44] and [45]. In contrast, our focus is on streamlining benchmarking practices for frequently used BEV datasets that lack established preprocessing and evaluation protocols. As such, our work is closely aligned with initiatives that define clear dataset partitions and evaluation metrics in the context of structured challenges such as [22], [23], and [24].
TABLE I: Time specifications of studies using BEV datasets.
@@ IV-B Dataset Splits
It is important to recognize that there is a notable imbalance across the maneuver classes within the highway datasets (the lane-change percentage is around $13\%$). To ensure that the scenarios are both dynamic and challenging, we adopt a stratified sampling approach, aiming to achieve a more balanced distribution of the TA maneuvers—the results of which are shown in Table III. We use this information to label the trajectories with the corresponding maneuver class, which can be used for assessing the ability of the model to predict specific maneuvers as well as intention prediction applications. Seven maneuver classes are employed based on the time-to-lane-change (TTLC), as described in Fig. 3.
@@ IV-C Coordinate System
Although the datasets are consistent in terms of the data they provide, some differences need to be addressed, such as the coordinate system used. This is mainly an issue when working with the highD dataset. To ensure that the samples are consistent across the datasets, the data are first divided into two separate groups based on the direction of travel. Next, the data are transformed to a common coordinate system where the $x$-axis points in the direction of the road, the $y$-axis points to the left, and the origin is moved to the lower left corner of the current road segment. For the other datasets, there are no such obvious issues, as all agents move in the same area. Regardless, we still place the origin in the approximate geographical center of the scene to ensure some consistency across recordings.
@@
A potential issue with using a global coordinate system is that it may introduce prediction bias, where the model learns scene-specific patterns rather than generalizable behavior—an essential property for zero-shot learning [46, 47]. While we leave this design choice to the user, one possible approach is to transform the scenario into a relative (agent-centric) coordinate system, such as the local frame of the TA.
IV-D Downsampling
@@
Most related works have used downsampling as a preprocessing step [29, 37, 34, 33, 35, 48, 38, 36, 39, 40, 41, 43], albeit with different integer factors. Although there are several potential computational benefits—allowing for faster prototyping and experimentation—it is important to consider the qualitative effect downsampling has on both the data and the model. Figure 4 is used to illustrate how different sampling rates can affect the quality of predicted trajectories using nominal driving conditions for the rounD dataset using a 1st-order method. In the figure, we see that for lower sampling rates, the simulated trajectories become overly jagged and unrealistic, which makes it difficult to assess the physical feasibility of the predicted trajectories. To ensure that the data are still representative of the original tracking data, we downsample it to a rate of $5$ Hz, corresponding to a sampling time of $0.2$ s. When studying the frequency spectrum of the tracking data, it was found that several signals have components above the new Nyquist frequency ($2.5$ Hz). Before downsampling, an anti-aliasing filter is applied to remove these high-frequency components using a 7th order Chebyshev type I filter (based on qualitative evaluations).
@@
Figure 4: Example simulation of a vehicle moving along a circular path with a radius of $16$ m, a speed of $50$ km/h (speed limit) for a duration of $5$ s using different sampling rates in a forward Euler integration scheme. The values represent nominal conditions for the rounD dataset. The figure is used to illustrate how different sampling rates can affect motion prediction quality. This is something that could be of interest when assessing the physical feasibility of the predicted trajectories.
IV-E Agent Features
IV-E1 Tracking Features
@@
Longitudinal and lateral accelerations are included as supplementary features. However, we discourage their direct use in the final model, as such measurements of high quality are rarely available in POV data. Estimated values (e.g., via finite differences) may still be used when appropriate.
@@
IV-E2 Agent Classes
The different datasets contain agents of different types, e.g., cars, bicycles, pedestrians, etc. We hypothesize that the model could benefit from the inclusion of this information, as different classes exhibit different behaviors. To allow for a more generalizable model, we propose a common set of agent classes across the datasets. The categories could be included as one-hot encoded features, or directly used as tokens in an embedding layer. The distribution of unique agents and their respective classes in the datasets is shown in Table V. We note that the number of unique agents in the INTERACTION dataset is not explicitly provided, although the accompanying paper reports a total of 40,054 recorded vehicles. For reference, our preprocessed data contains 676,931 vehicle trajectories, along with 32,722 trajectories corresponding to a single VRU class.
@@ IV-E2 Agent Classes
Figure 5: Example lane graphs for the rounD, inD, SIND, and INTERACTION datasets. The lane graphs are constructed using the available Lanelet files, which contain semantic information about the scene. The lane graphs are constructed by extracting lane boundaries and markings, coloring them according to their type, and connecting them with edges. Although these examples are simple, combined with the available traffic data, they could be sufficient for many trajectory prediction tasks. Using the provided functionality, users can develop more complex lane graphs based on the desired level of detail.
IV-F High-Definition Maps
@@
All but the highD dataset are accompanied by Lanelet 2 [28] files that can be used to extract semantic information about the scenes. Our toolbox includes functionality for basic lane-graph construction, designed to serve as a foundation that users can modify according to their specific needs and preferences. The resulting lane graph can be easily linked to agents in the scene, e.g., via a nearest-neighbor criterion. While the highD dataset lacks Lanelet files, available lane marking information are used to construct a simplified lane graph. Examples are shown in Fig. 5.
@@
IV-G Data Structure
V Evaluation Metrics
This section outlines the most commonly used metrics from the literature to evaluate motion prediction models, which are incorporated in the toolbox. The metrics are presented here for a single agent but can easily be extended to multi-agent problems by averaging over all scored agents in the scenario. Here, $\bm{x}_{k}\in\mathbb{R}^{2}$ refers to the ground truth position at time step $k$, $\hat{\bm{x}}_{k}$ refers to the respective prediction, and $N$ is the prediction horizon steps.
@@ V Evaluation Metrics
VI Future Directions
The introduction of the toolbox for the development of trajectory-prediction models for BEV datasets opens up several directions for future research. Here we outline some potential areas for further exploration.
@@ VI Future Directions
  • Zero-Shot Generalization: Explore model generalization and zero-shot learning across datasets to investigate how models can be applied or adapted across different scenarios.
  • Transfer Learning: Explore the potential of applying models, initially trained on one or more BEV datasets with subsequent fine-tuning, to other datasets within the suite or even to external datasets such as NuScenes, Argoverse 2, or WOMD. An example of how transfer learning can be applied to the INTERACTION and inD datasets using a simple encoder-decoder model is shown in Fig. 6.
Figure 6: Impact of transfer learning on validation performance for the inD and INTERACTION datasets. Fine-tuning after pretraining on the other dataset yields lower initial error and improved performance throughout training. These results highlight the potential of transfer learning for improved generalization and motivate further research in this direction.
VII Conclusions
This paper has presented a set of practices for developing trajectory prediction models using BEV datasets. The proposals have been motivated by analyzing the most common practices in the field. The methods have been implemented in an open-source toolbox designed to standardize preprocessing and performance evaluation procedures—thereby simplifying comparative studies. By alleviating the technical burden of developing research pipelines, the toolbox augments the potential for exploration and discoveries in the field of trajectory prediction.